The sum of twice a first number and three times a second number is 71. If the second number is subtracted from three times the first number the result is 2. Find the numbers.
Answer by erica65404(394)

2x+3y=71
3x-y=2
-y=-3x+2
y=3x-2
plug into equation
2x+3(3x-2)=71
2x+9x-6=71
11x=77
x=11
plug into y=3x-2
y=3(11)-2
y=31
The first number is 11 and the second number is 31